hello i need help on this problem 19c16? thanks
Answer by jim_thompson5910(35256)

n C r= (n!)/(r!(n-r)!)

19 C 16 = (19!)/(16!*(19-16)!)

19 C 16 = (19!)/(16!*3!)

19 C 16 = (19*18*17*16!)/(16!*3!)

19 C 16 = (19*18*17)/(3!)

19 C 16 = (19*18*17)/(3*2*1)

19 C 16 = (5814)/(6)

19 C 16 = 969

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------